Let 21 - 16sin^2(x) + 8cosx / 16cos^2(x) - 29 - 8√15sinx = 2. Find the largest value of 7cosx.
1
Expert's answer
2020-11-09T19:56:54-0500
"\dfrac{21-16\sin^2(x)+8\cos(x)}{16\cos^2(x)-29-8\sqrt{15}\sin(x)}=2"

Let "v=\cos(x), u=\sin(x)." Then


"u^2+v^2=1"

"16v^2-29-8\sqrt{15}u\not=0"

"21-16u^2+8v-32v^2+58+16\sqrt{15}u=0"

"(16v^2+8v+1)+2(16u^2+8\sqrt{15}u+15)=0"

"(4v+1)^2+2(4u+\sqrt{15})^2=0"

Then


"\begin{cases}
 4v+1=0 \\
 4u+\sqrt{15}=0
\end{cases}=>\begin{cases}
 v=-\dfrac{1}{4} \\
 u=-\dfrac{\sqrt{15}}{4}
\end{cases}"

Check


"u^2+v^2=(-\dfrac{1}{4})^2+(-\dfrac{\sqrt{15}}{4})^2=1, True"

"16v^2-29-8\sqrt{15}u="

"=16(-\dfrac{1}{4})^2-29-8\sqrt{15}(-\dfrac{\sqrt{15}}{4})=1-29+30="

"=2\not=0, True"

Hence

"\cos(x)=v=-\dfrac{1}{4}"

"7\cos(x)=-\dfrac{7}{4}=-1.75"